Indicate whether or not each of the following stru
Answer #1

Huckel’s Rule for Aromaticity is defined as,

Any Cyclic, highly conjugated, planar structure containing [4n+2] pi electrons (n = whole number 0,1,2………. And not the ring number) is Aromatic in nature and has extraordinary stability.

In (A) 1 C is sp3 hybridized and hence planarity of ring destroyed and cyclic conjugation also not possible hence (A) is Non-aromatic in nature.

In (B) –It’s a cyclic 5 C system 4 C which are doubly bonded are sp2 hybridized (4 pi electrons) and 1 C a carbanion which though sp3 hybridised but as conjugates its lone pair of electron in pi electron system gets planarity. Also 4 pi electrons + 2 electrons in –ve charge on carbanion makes total of 6 electrons i.e [4n+2] electrons (n = 1 makes it =4 x1 + 2 = 6 pi electrons). Hence The cyclopentadinyl anion is a Cyclic, highly conjugated, Planar structure with [4n+2] pi electrons system and thus is Aromatic by Huckel’s rule.